(7z^2 + 7z/ 4z +8) x (z^2 -4/ z^3 +2z^2 + z)
A) 7(z-2)/4(z+1)
B) 7(z+2)/4(z-1)
C) 7z (z+1)/4(z+2)
D) 7z (z-1)/4z(z+2)

Answer by richwmiller(17219)

factor
7z^2+7z=(7z(z+1)
z^3+2z^2 + z)=z *(z+1)^2
4z +8=4*(z+2)
z^2 -4=(z+2)*(z-2)
so we have
(7z(z+1)/ 4*(z+2) * (z+2)*(z-2)/z *(z+1)^2
remove z/z
(7(z+1)/ 4*(z+2) * (z+2)*(z-2)/(z+1)^2
remove (z+2) /(z+2)
(7(z+1)/ 4 * (z-2)/ (z+1)^2
remove (z+1) /(z+1)
7/ 4 * (z-2)/(z+1)
combine
7(z-2)/4*(z+1)
A) 7(z-2)/4(z+1)